40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 12 11:53:29 2018, aaron, Update, General, Power Outage IMG_0033.HEICIMG_1027.HEICIMG_2605.HEIC
    Reply  Thu Dec 13 01:26:34 2018, gautam, Update, General, Power Outage recovery 
       Reply  Thu Dec 13 12:06:35 2018, gautam, Update, General, Power Outage recovery 
          Reply  Thu Dec 13 20:10:08 2018, Koji, Update, General, Power Outage recovery OMRON_S82K-00712.JPGreset_button_repaired1.JPGreset_button_repaired2.JPGgray_box.JPGgray_box2.JPG
Message ID: 14353     Entry time: Thu Dec 13 20:10:08 2018     In reply to: 14351
Author: Koji 
Type: Update 
Category: General 
Subject: Power Outage recovery 

[Gautam, Aaron, Koji]

The PSL interlock system was fixed and now the 40m lab is laser hazard as usual.


- The schematic diagram of the interlock system D1200192
- We have opened the interlock box. Immediately we found that the DC switching supply (OMRON S82K-00712) is not functioning anymore.  (Attachment #1)
- We could not remove the module as the power supply was attached on the DIN rail. We decided to leave the broken supply there (it is still AC powered with no DC output).

- Instead, we brought a DC supply adapter from somewhere and chopped the head so that we can hook it up on the crimping-type quick connects. In Attachment #1, the gray is +12V, and the orange and black lines are GND.

- Upon the inspection, the wires of the "door interlock reset button" fell off and the momentary switch (GRAYHILL 30-05-01-502-03) got broken. So it was replaced with another momentary swicth, which is way smaller than the original unfortunately. (Attachments 2 and 3)

- Once the DC supply adapter was pluged to an AC tap, we heard the sounds of the relays working, and we recovered the laser hazard lamps, PSL door alerm lamps. Also it was confirmed that the PSL innolight is operatable now. 

- BTW, there is the big switch box on the wall close to the PSL enclosure. Some of the green lamps were gone. We found that we have plenty of spare lamps and relays inside of the box. So we replaced the bulbs and know the A.C. lights are functioning. (Attachments 4 & 5)

Attachment 1: OMRON_S82K-00712.JPG  706 kB  | Hide | Hide all
OMRON_S82K-00712.JPG
Attachment 2: reset_button_repaired1.JPG  376 kB  | Hide | Hide all
reset_button_repaired1.JPG
Attachment 3: reset_button_repaired2.JPG  308 kB  | Hide | Hide all
reset_button_repaired2.JPG
Attachment 4: gray_box.JPG  315 kB  | Hide | Hide all
gray_box.JPG
Attachment 5: gray_box2.JPG  602 kB  | Hide | Hide all
gray_box2.JPG
ELOG V3.1.3-